Details.

RR intalled on DX

08:30am

Talked with Guido reason use of DX Adsec. Guido said OK to use the Adsec, if we have any failures, call him.

08:40 am

Briefing talk.

Refine the pupils for Bin1, and acquire IM with the RR..

09:05am

WFS power oN
  • OK
Ocam
  • OK
Hexapod positions:
  • Initial guess(September 2019)
    • 0->-0.68 0.975 1.6 -399 -288
    • 1-> removed coma: 1.183 1.2545 1.62 -428.1439 93.8293
    • 2-> 1.18 1.25 1.73 -446 -64 (Sep18,2019)
    • 3->1.41 -1.75 1.2 -157.6 -22.7 (Jan 2020)

Parallel work during the acquition of the source.

09:30 Doug & Enrico Refining pupil of Bin1

ao>refine_pup, '20200131_180431',1

We want to add a full circle of SA at the external edge and then verify their illumination with the IM RMS Check

We did it with refine_pup.pro, and created NEW ENLARGED BIN1 PUP: 20221201_171333

Loaded on the WFS and working. They have 42 pix on the diameter


9:35am
  • Released telescope. Green light from the telescope manager.
  • Luci Authorise.
Load the Jan2020 hexapod postions.
  • Not light on th CCD47.
  • Scan various postions moving the Rx and Ry.
  • No luck.
Go to hexa_spiral(500e-6,15,/apply)

Found a possible source in the CCD47, it was not too bright

Not clear theat the spot is the source. Expected to see two bright spots that change in focus.

launch Luci wide field to see where it is the source.
  • Filter J, 3 sec.
  • Set luci in live mode
  • Luci image shows the center of the RR and a source.
  • Try to align the source to the center.
  • Better to have the center RR in the center.
  • Bring the spot to the center of the RR obscurtation. Ths is a good technique. Defocused PSF and observed two full pupls (real and the second reflection)
  • Improve the alignment moving the stage and offload with the Rx and Ry. Using this technique was able to have a point source centerred in the CCD47
Hexapod: -0.88, 0.43,-0.66,-248.1,-41,0

11:15 AM

good spot identified on LUCI2

We loose time trying to aling the wrong reflection on the LUCI dichroic.

12:15 PM

Cube had a bad rotation by 90 deg!

Rotating by 90 we have btighter source and with no vignetting on LUCI. Open loopTN To record position of M2-M3 Cube TN 20221201_191742

12:30PM

Close loop with yesterday Rec.

TN:_193343

Acq Inmat using the RR@900Hz
  • check phtons in ocam min 500 threshold.
Acquire IM
  • KL_29
  • 20220607_184919
  • close loop
  • iter=4
  • Rec:20221130_192802
  • gain:0.3,0.05,0.05
  • WFS setup:
  • bin1
  • 600Hz,3l/D
  • EMG:67
  • output IM:20221201_194333
Analyse.
  • observed some dent in the mode tip
  • Reconstruct 500.
  • reconstructo modes (insert plot)
  • observed the ring that it related to the edge
  • Reconstructor: 20221201_194742.
12:59PM

Test Close loop
  • no dist:
  • 20221201_195928,_200010
Analyse: IM:20221201_194333
  • cut 629
  • add the list of bad modes: 504, 540, 541, 574, 575, 576, 600
  • REC:20221201_200100
01:07PM

test Close loop.

TN:20221201_200721
  • Skip frames 100%.
  • Pupil jumping around.
  • large vibratons.
  • loop open
  • too much vibration.
1:15PM

New Rec
  • to improve the edge control and save a flat.
  • Rec:20221201_200100
  • gainv:0.3,0.05,0.05
  • iter=4
  • wfs
  • 600Hz
  • 3l/D
  • engain:63
  • close loop
  • output file:20221201_201536
  • raw signal plot
    • many below 0.8
Analyse:
  • cut 629
  • list of modes to cut:504, 540, 541, 574, 575, 576, 600
  • outout Rec:20221201_201809
  • still in the Rec modes the edge.
  • it looks not too good.
test close loop.
  • gain TT0.85,0.05,0.05
  • TN:202816
  • stable.
  • skip frame.
  • vibrations telescope/RR,swing arm,etc..
  • loop open

  • Observed Actuator224 with a large force; current:-0.22
  • save flat: 20221201_600m_2.sav
  • apply flat.
  • TN:_203620
  • save flats 20221201_600m_3..4.sav
  • close loop
  • gain 1, 0.05,0.05
  • open
  • load flat: 20221201_600m_5
  • load flat; act 228 triggerd in forces.
  • 402 modes applied.
01:41PM

load yesterday flat

20221130_0930-10lD
  • Stop AO didn't load the flat, the adsec remains in almost rest postions.
  • loaded the flat it was in half signal. Still some forces.(?) to talk with Guido

Acquire Inmat using a trick for open loop and using other pupils.

Method: open loop fast change bin2 to 1 from the wfscontrol loop and start to acquire the Inmat in open loop

During the change of binning the ocam shows to have a low photons, check filter 1, home filter and setup again from the wfscontrol gui; bin1 600Hz, emgain=67 3l/d.
  • try another Bin2 to Bin1
  • bin1
  • emgain:67
  • OCAm magnitude mag~6
2:00PM

Stop

Moving to the IIR discussion and OCAM.

Power off WFS, Adsec flat.

-- JuanCarlosGuerra - 01 Dec 2022
Topic revision: r4 - 02 Dec 2022, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback